A farmer grows only pumpkins and corn on her 420 -acre farm. This year she wants to plant 250 more acres of corn than pumpkins. How many acres of each crop does the farmer need to plant?
Answer by ikleyn(52781)

.
A farmer grows only pumpkins and corn on her 420 -acre farm.
This year she wants to plant 250 more acres of corn than pumpkins.
How many acres of each crop does the farmer need to plant?
~~~~~~~~~~~~~~~~~~~~~~~

Let P be the pumpkins area (in acres).

Then the corn area is (P+250) acres, according to the condition.



With it, write the total area equation as you read the text

    P + (P+250) = 420  acres.


Simplify and find P

    2P = 420 - 250 = 170

     P             = 170/2 = 85.


ANSWER.  85 acres of pumpkins and 85 + 250 = 335 acres of corn.


CHECK.   The total area is  85 + 335 = 420 acres.    ! Correct !


Solved.


                There is another way to solve the problem.


You can write the system of two equations

    C + P = 420

    C - P = 250.


To solve it, add the equations.


Doing this way to the end, you will get the same answer.


Solved   (in two ways).


Happy learning  (!)